# Explicit Formula of the Infinite Energy Solutions for the Exterior 2D Div-Curl Problem with Dirichlet Boundary Condition

### v2

@@ 5 Numerical results
Since the no-slip condition (30) is fullfiled, then there exists a stream function $\psi$, such that
@@
The condition (30) is a particular case of slip interaction between solid and fluid, which is expressed by the formula:
@@
The equation (29) goes to Poisson equation under curl operator:
@@ 6 Discussion
The velocity field vanishes on the boundary in some approximation (Figures 1, 3) since in the computational process only a finite set of relations (27) can be imposed ($k<20$ in the presented experiment), and the boundary velocity may produce high-frequency fluctuations with small amplitude corresponding to non-zero Fourier coefficients for large indexes $k$.
Regarding the geometry of the exterior domain $\Omega$, in the current research its boundary consists of two connected components (including point at infinity). But the presented approach can be extended to the case of more than two connected components, when the fluid flows through several solids. If these solids are far apart from each other, then (27) becomes an approximation of the no-slip condition for each object.
In exterior of the line segment the statement of the div-curl problem requires some explanation. In that case the line segment has "two sides" in some sense, and no-slip condition must be imposed on both sides of it. It can be considered as a limiting case of ellipses when its thickness goes to zero. Another way the definition of the solution can be obtained via Riemann mapping of the div-curl problem into the exterior of the disk, where we can correctly define the solution of the equation.
@@ 7 Conclusion
Integral constraints (27) imposed on vorticity function $w$ and divergence $\rho$ give unique solvability of the exterior div-curl problem with no-slip boundary condition around solid. It can be solved numerically by reducing to Poisson problem. This approach can be extended to flows around several streamlined solids.
For solenoidal flows no-slip condition generates a series of integral constraints on vorticity (orthogonality relations). Under these restrictions the div-curl problem can be solved in correct formulation with more typical boundary conditions - slip condition combined with zero circulation. The source of vorticity data, which satisfies such integral constraints, can be derived by solving any equations of motion written in vorticity form that describe fluid dynamics - Stokes, Navier-Stokes, and Boussinesq equations. And the div-curl problem becomes the crucial step of the simulation procedure in the planar computational fluid dynamics.

### v3

@@ 5 Numerical results
Since the no-slip condition (30) is fullfiled, then there exists a stream function $\psi$
@@
The condition (30) is a particular case of a slip interaction between solid and fluid, which is expressed by the formula:
@@
Applying the curl operator to (29) it goes to Poisson equation:
@@ 6 Discussion
The velocity field vanishes on the boundary in some approximation (Figures 1, 3) since in the computational process only a finite set of relations (27) can be imposed ($k<20$ in the presented experiment), and the boundary velocity may produce high-frequency fluctuations with small amplitude corresponding to non-zero Fourier coefficients for large indices $k$.
Regarding the geometry of the exterior domain $\Omega$, in the current research its boundary consists of two connected components (including point at infinity). But the presented approach can be extended to the case of more than two connected components, when the fluid flows through several solids.
In exterior of the line segment the statement of the div-curl problem has its own peculiarities. In that case the line segment has "two sides" in some sense, and no-slip condition must be imposed on both sides of it. It can be considered as a limiting case of ellipses when its thickness goes to zero. Another way the definition of the solution can be obtained via Riemann mapping of the div-curl problem into the exterior of the disk, where we can correctly define the solution of the equation.
@@ 7 Conclusion
Integral constraints (27) imposed on vorticity function $w$, divergence $\rho$ and boundary function $\mathbf{g}$ give unique solvability of the exterior div-curl. It can be solved numerically by reducing to Poisson problem. This approach can be extended to flows around several streamlined solids.
For solenoidal flows no-slip condition generates a series of integral constraints on vorticity (orthogonality relations). Under these restrictions the div-curl problem can be numerically solved with more typical boundary conditions - slip condition combined with zero circulation. The source of vorticity data, which satisfies such integral constraints, can be derived by solving any equations of motion written in vorticity form that describe fluid dynamics - Stokes, Navier-Stokes, and Boussinesq equations. And the div-curl problem becomes the crucial step of the simulation procedure in the planar computational fluid dynamics.